# Semiparametric Skew-Elliptical Distributions For High-Dimensional Graphical Models

### v1

@@ 5.1 Numerical Simulations
Figure 1: Comparison of average ROC curves, using the Glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 100 trials. We consider, by column, three different data generations. And, by row, three different dataset sizes.
@@ 5.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimator and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the Normal distribution (Cont, (2001)).
@@
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events thus a huge presence of outlier value during the analyzed period, which lead to a departure from the assumption of Normal distribution for log-returns.
@@
To estimate the graph, we determine the tuning parameter $\lambda$ using Extended Bayesian Information Criteria (EBIC) developed in Chen and Chen, (2008):
$\operatorname{EBIC}=-2L+Elog(n)+4\gamma Elog(p).$ (23)
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$. The tuning parameter is selected as $\lambda_{G}=0.54$ since it minimizes the scores of the first-order differences in EBIC.
@@
For estimating the undirected graph conforming to the SUT distribution, using the sut Kendall estimator, we need to estimate the degrees of freedom of the log-returns distributions. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the $R$ package sn developed by Azzalini, (2023).
@@
Figure 3 shows the estimated graphs using npn Kendall, csn Kendall, and sut Kendall estimators.
Figure 3: The resulting graphs using the (a) csn Kendall estimator, (b) sut Kendall estimator, and (c) npn Kendall estimator for S&P 500 stock data from $04/01/2016$ to $31/12/2022$.
@@
The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks). The stocks that belong to the same sector tend to cluster. Nevertheless, csn Kendall and sut Kendall show in better way the grouping of the stocks, thus improving the interpretability of the visualization.
@@
Stocks at the periphery of the network are uncorrelated to the others, during the considered period, thus they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) shows that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in high-dimension, a network approach can be valuable for identifying stocks with the least correlation to the market.
@@
Table 4 shows that introducing the skewness and kurtosis components results in a sparser network estimation compared to the Normal case, thus the elliptical skew-(S)KEPTIC estimator improves the interpretability of the graph.
@@
Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation. SK = SKEPTIC, SNK = elliptical skew-SKEPTIC, STK = elliptical skew-KEPTIC. DIM = dimension of the dataset. The notation ”$>$” indicates that one method has a greater number of edges compared to the other.
@@ 6 Concluding remarks
We propose a rank-based correlation matrix estimator, named the elliptical skew-(S)KEPTIC estimator which extends to the class of the skew elliptical distribution in estimating the high-dimensional graphical models. The elliptical skew-(S)KEPTIC estimator results for Kendall’s tau in the class of skew-elliptical distribution, and it also encompasses Spearman’s rho when the class of distribution reduces to the skew-Normal. We expand the SKEPTIC estimator in Liu et al., 2012a to the class of semiparametric meta skew-elliptical distribution introduced throughout the paper.
Our analysis is not asymptotic, and the estimator is defined in a way in which we do not explicitly estimate the marginal transformation of the data. Theoretically, the slight reduction in the robustness of the estimator is compensated by the extension of the class of distributions over which we estimate the graphical model.
We illustrate a method for modeling data that conforms to a $\operatorname{MS}\mathcal{E}$ distribution by treating the data generation process as a mixture of Normal and half-Normal distributions. For both the multivariate random variables, the SKEPTIC estimator is defined. Thus, the estimator is a sum of two SKEPTIC estimators and a matrix accounting for skewness. This extension yields a sparser pattern compared to a Gaussian graphical model, resulting in improved interpretability of the resulting graphs. Further, the elliptical skew-(S)KEPTIC estimator shows a slower exponential concentration in $\|\cdot\|_{\max}$ rate than the SKEPTIC estimator. Nevertheless, our findings show a reliable outcome that is useful when dealing with data that exhibit high skewness when it cannot be neglected. Our approach’s effectiveness is validated through optimal graph recoveries in numerical simulations.
@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
Let $\hat{\rho}_{ij}$ denote the Spearman’s rho estimator of $\Sigma^{0}_{ij}=Corr(X_{i},X_{j})$, the estimator is biased (Zimmerman et al., , 2003).
$\mathbb{E}(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right],$ (24)
@@
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\text{where: }\quad a_{ij}=\frac{\pi\mathbb{E}(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$ (25)
The term $|a_{ij}|\leq\frac{\pi}{n-2}$, therefore for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$ and thus:
$\mathbb{P}\left(\left|\hat{S}^{\rho,\alpha}_{ij}-\Sigma^{0}_{ij}\right|>t\right)=$ (26)
$=p^{2}\mathbb{P}\left(\left|4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-2\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(2\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=p^{2}\mathbb{P}\left(\left|2\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)-\sin\left(\frac{\pi}{6}\mathbb{E}(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})+\frac{6}{\pi}a_{ij}\right|>\frac{3}{\pi}t\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\left|\frac{6}{\pi}a_{ij}\right|\right)$
$\leq p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{3}{\pi}t-\frac{1}{\pi}t\right)$
$=p^{2}\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$
The correlation matrix using Spearman rho, $\hat{\Sigma}_{\rho}$, can be decomposed, using modified Cholesky decomposition as $\hat{\Sigma}_{\rho}=\hat{L}\hat{D}_{\kappa}\hat{L}^{\top}$, where $\hat{D}_{\kappa}=\operatorname{I}_{p}$. In our case, we have: $A_{\hat{\rho}_{ij}}=\hat{L}(I-\hat{D}^{2})\hat{L}^{\top}$. By definition the $k$-th element of $(1-\hat{d}_{kk}^{2})<1$. Thus:
$\mathbb{P}\left(\left|2\left(A_{\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\mathbb{P}\left(\left|2\left({\hat{\rho}_{ij}}\right)+\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)$ (27)
We prove the bound of the estimator by rewriting Spearman’s rho in terms of the U-statistic and using the Hoeffding inequality. This implies:
$\mathbb{P}\left(\sup_{ij}\left|\hat{S}^{\rho,\alpha}-\mathbb{E}(\hat{\rho}_{ij})\right|>\frac{2}{\pi}t\right)\leq\exp\left(-\frac{nt^{2}}{16\pi^{2}}\right)$ (28)
A.2 Proof of Theorem 4.2

### v2

@@ 6.1 Numerical Simulations
Figure 1: Comparison of average ROC curves, using the glasso algorithm, of the elliptical skew-(S)KEPTIC estimators with SKEPTIC estimators. Results are determined over 50 trials. We consider, by row, three different data generations and, by column, three different dataset sizes.
Table 2: Quantitative comparisons of the 5 correlation matrix estimators using different distributions to generate data, and compute the $\Omega$ matrix. The graphs are estimated using the glasso algorithm. The values are in percentages.
SKEPTIC ($\hat{\rho}$) SKEPTIC ($\hat{\tau}$) SKEW-SKEPTIC ($\hat{\rho}$) SKEW-SKEPTIC ($\hat{\tau}$) PEARSON
$n$ $p$ AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR AUC FPR FNR
N DATA 200 100 85.0 2.06 27.4 85.0 2.07 27.5 92.0 11.7 4.8 92.0 12.8 3.7 86.0 1.8 25.7
s.e (10.0) (2.6) (21.6) (10.0) (2.6) (21.6) (2.0) (7.7) (3.2) (3.0) (7.8) (2.2) (10.0) (2.3) (21.7)
1000 250 98.3 2.7 0.7 98.0 2.7 0.7 93.0 13.2 6.3 90.5 16.7 11.8 98.5 2.4 0.6
s.e (3.0) (5.5) (1.3) (3.0) (5.5) (1.4) (12.5) (24.6) (24.3) (16.0) (31.4) (32.3) (2.4) (4.9) (1.1)
500 300 94.0 1.3 9.8 94.0 1.3 9.7 96.0 7.8 0.3 96.0 8.8 0.1 95.0 1.1 9.4
s.e (7.0) (2.4) (14.4) (7.0) (2.5) (14.2) (4.0) (8.3) (0.5) (4.0) (8.7) (0.2) (7.0) (2.1) (14.1)
CSN DATA 200 100 86.0 5.0 22.8 86.0 5.0 22.2 88.0 17.3 6.3 88.0 18.5 5.7 83.0 4.4 28.9
s.e (10.0) (4.9) (21.1) (10.0) (4.9) (20.8) (4.0) (10.2) (2.1) (5.0) (10.2) (2.1) (10.0) (4.7) (24.1)
1000 250 98.0 1.6 3.2 98.0 1.6 3.0 96.0 8.3 0.7 94.2 11.0 4.5 96.0 1.3 6.7
s.e (3.0) (2.5) (6.6) (3.0) (2.5) (6.2) (5.7) (11.3) (8.1) (10.0) (20.6) (20.7) (6.0) (2.2) (12.5)
500 300 92.0 1.5 14.7 92.1 1.6 14.2 95.0 8.3 1.7 95.0 9.4 0.9 89.5 1.2 19.7
s.e (9.1) (2.2) (19.3) (8.9) (2.2) (18.8) (3.4) (7.8) (1.6) (4.0) (8.3) (0.6) (11.7) (1.8) (24.4)
SUT DATA 200 100 80.0 2.9 37.4 80.0 3.0 36.7 (-) (-) (-) 89.0 15.6 7.1 71.0 2.2 54.9
s.e (10.0) (3.8) (24.2) (10.0) (3.9) (37.4) (-) (-) (-) (2.0) (6.3) (1.7) (12.0) (3.2) (26.8)
1000 250 97.0 2.1 4.7 97.0 2.1 4.5 (-) (-) (-) 96.0 7.7 0.1 89.0 1.7 20.3
s.e (4.0) (4.5) (8.6) (4.0) (4.5) (8.3) (-) (-) (-) (3.7) (7.4) (0.01) (12.0) (4.2) (25.2)
500 300 89.3 1.9 19.8 89.6 1.9 18.9 (-) (-) (-) 94.4 9.7 1.5 80.0 1.6 37.6
s.e (9.1) (3.7) (19.8) (9.0) (3.7) (19.5) (-) (-) (-) (3.3) (6.9) (0.7) (14.4) (3.3) (30.5)
AUC = Area Under Curve, FPR = False Positive Rate, and FNR = False Negative Rate.
@@ 6.2 Empirical Analysis: S&P500 Log-Returns
In the empirical analysis, we study the conditional dependence relations on daily returns of S&P500500 companies in the period from $04/01/2016$ to $31/12/2022$ downloaded from Yahoo! Finance, using the proposed elliptical skew-(S)KEPTIC estimators and comparing it with the SKEPTIC. After cleaning the data, the dataset consists of 454 stocks, each counting 1762 observations. The log-returns of the stocks in the S&P500 index are empirically skewed and present heavier tails than the normal distribution (Cont, (2001)).
@@
As a motivating example, we benchmark the analysis of Han and Liu, (2014), reporting the asymmetry of log-returns in Figure 2, where we plot histograms of the returns of three stocks in our sample ”eBay Inc.”, ”Coca-Cola” and ”Wells Fargo” belonging to three different sectors, respectively, Consumer Discretionary, Consumer Staples, and Financials. We appreciate the departure from normality reflected in leptokurtic distribution with a relevant skewness component. Moreover, we investigate the Gaussian assumption of the asset log-returns by computing normality tests with different significance levels at 5% and 1%. Table 3 indicates a clear and complete deviation from normality in the stock data. This could be attributed to the presence of exceptional events, thus a huge presence of outlier values during the analyzed period, which lead to a departure from the assumption of normal distribution for log-returns.
@@
To estimate the graphs, given the large number of variables relative to observations we determine the tuning parameter $\lambda_{G}$ using the Extended Bayesian Information Criteria (EBIC) of Chen and Chen, (2008) to ensure a balanced trade-off between model fit and complexity.
$\operatorname{EBIC}=-2L+E\log(n)+4\gamma E\log(p).$ (19)
@@
The hyperparameter $\gamma\in\left[0,1\right]$, with greater values of $\gamma$ implying a stronger penalization, $L$ denotes the log-likelihood, $n$ the sample size, $E$ the number of non-zero edges and $p$ the number of nodes. We set $\gamma=1$, ensuring a strong penalization, as desired. The tuning parameter $\lambda_{G}$ is set equal to $0.54$ the value that minimizes the scores of the first-order differences in EBIC. To align the log-returns to the unified skew-$t$ distribution for computing the elliptical skew-KEPTIC estimator, we estimate the degrees of freedom of the distribution. We follow Arellano-Valle and Genton, 2010a using the maximum likelihood estimation method. In likelihood methods, the underlying assumption is that the observations are independent, therefore, we apply the best ARIMA model to the series using the R function auto.arima to filter our returns into a white noise series, and we apply the likelihood method to the filtered series to estimate the degrees of freedom using the command mst.mle from the R package sn developed by Azzalini, (2023).
@@
Figure 3 shows the estimated graphs using SKEPTIC, elliptical skew-SKEPTIC, and elliptical skew-KEPITC estimators; all the correlation matrix estimators are based on Kendall’s tau statistic.
Figure 3: The resulting graphs of the log-returns of the S&P 500 stocks from $04/01/2016$ to $31/12/2022$. Kendall’s tau statistic is employed to estimate correlation. In the graph: (a) SKEPTIC estimator, (b) elliptical skew-SKEPTIC estimator, and (c) elliptical skew-KEPTIC estimator.
@@
The 454 stocks are categorized into 11 Global Industry Classification Standard (GICS) sectors, including Consumer Discretionary (51 stocks), Consumer Staples (33 stocks), Energy (21 stocks), Financials (67 stocks), Health Care (56 stocks), Industrials (69 stocks), Information Technology (59 stocks) Telecommunications Services (17 stocks), Materials (25 stocks), Real Estate (28 stocks), and Utilities (28 stocks).
@@
The stocks that belong to the same sector tend to cluster, and those at the periphery of the network are uncorrelated with the others during the considered period, thus, they are uncorrelated with the movement of the market. Peralta and Zareei, (2016) show that those stocks have a greater weight in constructing a portfolio using the Markowitz approach. Therefore, in large-dimensions, a network approach can be valuable for identifying stocks with the least correlation to the market.
@@
Table 4 demonstrates that incorporating skewness and kurtosis components leads to a sparser network estimation compared to the normal case.
@@
Table 4: Summary statistics of the estimated networks using Kendall’s tau statistic for the correlation matrix estimation. SK == SKEPTIC, SNK == elliptical skew-SKEPTIC, STK == elliptical skew-KEPTIC. DIM = dimension of the dataset. The notation ”>>” indicates that one method has a greater number of edges compared to the other.
@@ 7 Concluding remarks
We propose rank-based correlation matrix estimators, named the elliptical skew-(S)KEPTIC estimators, to provide semiparametric estimation of graphical models built on meta skew-elliptical distributions. The proposed framework includes two classes of estimators: the elliptical skew-KEPTIC estimator which employs Kendall’s tau statistic, and the elliptical skew-SKEPTIC estimator which also uses Spearman’s rho statistic to estimate the correlation matrix.
We study the theoretical properties of the rank-based correlation matrix estimators and establish concentration bounds. These bounds yield statistical rates of convergence under the elementwise maximum norm for both the correlation matrix estimation and the resulting estimation of the precision matrix. We also study the sparsistency of the precision matrix, evaluating the performance of the elliptical skew-(S)KEPTIC estimators in correctly recovering the edges of the graphical model.
The empirical analysis and numerical simulations demonstrate that the elliptical skew-(S)KEPTIC estimators yield the optimal performance when the data deviates from normality. Applied to real data, specifically the log-returns of S&P500 constituent stocks, the resulting graphs are sparser compared to those estimated under normality.
Our findings show a reliable framework that is useful when dealing with data that exhibit high skewness which cannot be neglected, and its effectiveness is validated through the optimal graph recoveries of the numerical simulations.
@@ Appendix A Proofs of elliptical skew-(S)KEPTIC estimators
We consider $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$, where $f(\boldsymbol{X})\sim V_{0}^{-1/2}Z_{0}$, and
$Z_{0}\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$ where $D_{\kappa}=I_{p}$, and $V_{0}=1$. We compute the correlation matrix of $\boldsymbol{X}$ using Kendall’s tau statistic, $\Sigma^{0}_{\tau}=\left[\tau_{ij}\right]$, and we decompose $\Sigma^{0}_{\tau}=L^{\top}D_{\kappa}L$, aligning it $\Sigma^{0}_{\tau}$ to the assumption of $\boldsymbol{X}\sim MS\mathcal{E}_{p,p}(\Omega^{-1},D_{\alpha}L,h^{(2p)},0,D_{\kappa};f)$ exploiting (8).
$\text{cov}(\boldsymbol{X})$ $=\left(L^{\top}(I-D^{2})L\right)^{-1},$
$=(L^{\top}ML)^{-1}\quad\text{where:}\quad M=I-D^{2}.$
The matrices $I$ and $D^{2}$ are both diagonal matrices, and the matrix $M$ is. Thus, we have a new decomposition of the correlation matrix in which the skewness parameter $\alpha$, belonging to the matrix $D_{\alpha}$, contributes to the determination of the linear relations. The matrix $(L^{\top}ML)$ is:
where the generic $m_{k,k}$ term is:
$m_{kk}=I_{kk}-D^{2}_{kk}=1-\frac{2}{\pi}\frac{\alpha_{kk}^{2}}{1+\alpha_{kk}^{2}}.$
Let $A_{{\tau}}\equiv L^{\top}ML$ denote the rescaled final correlation matrix of $\boldsymbol{X}$, estimated using Kendall’s tau statistic. We recall that the stochastic representation (7) of the CSN is a sum of two comonotic copula functions that have the invariance property, so the SKEPTIC estimator is the same for both the distributions, thus the elliptical skew-KEPTIC estimator is the sum of two SKEPTIC estimators and an estimated matrix, $A_{\tau}$, from $f(\boldsymbol{X})\sim\operatorname{CSN}_{p,p}(0,\Omega^{-1},D_{\alpha}L,0,D_{\kappa})$.
When $V_{0}\sim Gamma(\nu/2,\nu/2)$, we have that $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Sigma,D_{\alpha}L,\nu,0,D_{\kappa})$ or equivalently: $f(\boldsymbol{X})\sim\frac{Z_{0}}{\sqrt{V_{0}}}$. Since the positive support of Gamma random variable, and the fact that Kendall’s tau is invariant under monotone transformations, we state that the skew-KEPTIC estimator maintains the same ordinal measures of dependencies of the closed skew-normal distribution even when $f(\boldsymbol{X})\sim\operatorname{SUT}_{p,p}(0,\Omega^{-1},D_{\alpha}L,\nu,0,D_{\kappa})$. ∎
A.2 Proof of Lemma 3.5
Proof.
We recall Lemma 2.1, where for the normal distribution, and for $i\neq j$, we have: ${S}_{ij}=2\sin\left(\frac{\pi}{6}{\rho}_{ij}\right)=\sin\left(\frac{\pi}{2}{\tau}_{ij}\right)$. Let $A_{\rho}$ denote the final correlation matrix of $\boldsymbol{X}$ with Spearman’s rho statistic, as in the normal distribution, we have: ${S}^{\rho,\alpha}_{ij}\equiv 4\sin\left(\frac{\pi}{6}A_{{\rho}_{t}{ij}}\right)=2\sin\left(\frac{\pi}{2}A_{{\tau}_{ij}}\right)$. ∎
A.3 Proof of Theorem 5.1
Proof.
We have: $\hat{S}^{\rho,\alpha}_{ij}=$ $4\sin\left(\frac{\pi}{6}A_{\hat{\rho}_{ij}}\right)$ and we introduce $\Sigma^{\rho}$, the correlation matrix using Spearman’s rho statistic. We know that $A_{\rho}=L^{\top}ML$, where $M=I-D^{2}$. By Definition: $\Sigma^{\rho}=L^{\top}D_{\kappa}L$. This implies:
$\left\lVert\widehat{D}_{\kappa}\right\rVert_{\max}\geq\left\lVert\widehat{M}\right\rVert_{\max},$
$\left\lVert L^{\top}\widehat{D}_{\kappa}L\right\rVert_{\max}\geq\left\lVert L^{\top}\widehat{M}L\right\rVert_{\max},$
$\left\lVert\widehat{\Sigma}^{\rho}\right\rVert_{\max}\geq\left\lVert A_{\widehat{\rho}}\right\rVert_{\max},$
$\left\lVert 4\sin\left(\frac{\pi}{6}\widehat{\Sigma}^{\rho}\right)\right\rVert_{\max}\geq\left\lVert 4\sin\left(\frac{\pi}{6}A_{\widehat{\rho}}\right)\right\rVert_{\max}.$
[…]
$E(\widehat{\rho}_{ij})=\frac{6}{\pi(n+1)}\left[\arcsin\left(\Sigma_{ij}^{0}\right)+(n-2)\arcsin\left(\frac{\Sigma_{ij}^{0}}{2}\right)\right].$ (20)
@@
$\Sigma^{0}_{ij}=2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right),\quad\text{where }a_{ij}=\frac{\pi E(\hat{\rho}_{ij})-2\arcsin(\Sigma^{0}_{ij})}{2(n-2)}.$
The term $|a_{ij}|\leq\frac{\pi}{n-2}$. Therefore, for all $n>\frac{6\pi}{t}+2$, we have $|a_{ij}|<\frac{t}{6}$, and thus:
$\text{pr}\left(\left\lVert 4\sin\left(\frac{\pi}{6}\hat{\Sigma}^{\rho}\right)-\Sigma^{0}\right\rVert_{\max}>t\right)$ $=\text{pr}\left(\max_{ij}\left|4\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-2\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>t\right)$
$=\text{pr}\left(\max_{ij}\left|2\sin\left(\frac{\pi}{6}\hat{\rho}_{ij}\right)-\sin\left(\frac{\pi}{6}E(\hat{\rho}_{ij})+a_{ij}\right)\right|>\frac{t}{2}\right)$
$\leq\text{pr}\left(\max_{ij}\frac{2\pi}{3}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})-a_{ij}\right|>\frac{t}{2}\right)\quad\text{(Lipshitz function)}$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-|a_{ij}|\right)$
$\leq\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{3t}{4\pi}-\frac{t}{6}\right)$
$=\text{pr}\left(\max_{ij}\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)$
$\leq\sum_{ij}\text{pr}\left(\left|\hat{\rho}_{ij}-E(\hat{\rho}_{ij})\right|>\frac{(9-2\pi)t}{(2\pi-3)}\right)\quad\text{(Boole's inequality)}$
$\leq p^{2}\exp\left(-\frac{nt^{2}(9-2\pi)^{2}}{(2\pi-3)^{2}}\right)\quad\text{(Hoeffding's inequality)}$
Setting:
$\beta=p^{2}\exp\left(\frac{(9-2\pi)^{2}nt^{2}}{(2\pi-3)^{2}}\right)$
and solving for $t$, we get:
$t=\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left(\frac{p^{2}}{\beta}\right)}$
When $\beta=\frac{1}{p}$, we have $\sqrt{\frac{(2\pi-3)^{2}}{(9-2\pi)^{2}n}\log\left({p^{3}}\right)}$. ∎
A.4 Proof of Theorem 5.2